OpenStudy (anonymous):

On a trip from Williamsburg, Virginia to Orlando, Florida, a family wants to travel at least 250 miles in the first 5 hours of driving. What should their average speed be in order to meet their goal? s<

OpenStudy (anonymous):

v = distance/time = 250/5 = 50 mph

OpenStudy (saifoo.khan):

\[Speed = \frac{distance}{time}\]